The larger of two numbers is 12 more than the smaller. The sum of the number is 36. What are the numbers? Write a equation.
Answer by checkley71(8403)

X+(X+12)=36
X+X+12=36
2X=36-12
2X=24
X=12
X=12=24
PROOF
12+24=36
36=36